# Towards a Deeper Understanding of Transformer for Residential Non-intrusive Load Monitoring

### v2

@@ Abstract
Smart grid, Non-intrusive Load Monitoring (NILM), Energy disaggregation, Transformer.I Introduction
@@ I-A Related works
Deep learning approach: Deep learning has emerged from solving mainstream computer vision problems and influenced other data-driven research fields. Kelly and Knottenbelt are one of the first to propose a deep learning based NILM framework [7] that outperformed classical machine learning methods. The advantage of deep learning architectures is that they can directly process the raw power signal without the need for manual feature extraction or data tagging. They also accommodate noise signals and unknown appliances without assuming that the aggregated power signal fully represents the sum of the exact individual appliance power signal. Zhang et al. [8] enhanced their CNN model’s capacity which surpassed the previous methods. However, CNN models effectively capture the local features within power consumption sequences but they may not adequately capture global features and dependency correlations between different positions in the sequence.
Attention mechanisms [9], originally formulated in the field of natural language processing, happen to show promising results with NILM problems, especially with transformers. Yue’s BERT4NILM [10] incorporates a bidirectional transformer architecture. Followed by their work methods such as, Sykiotis’s ELECTRIcity-NILM [11], Wang’s Midformer [12], Yue’s ELTransformer[13], Kamyshev’s COLD[14], Zhou’s TTRNet[15] employed transformer in NILM. Transformer’s capability to capture long-range temporal dependencies made them highly applicable for time-series power consumption data.
@@ II Preliminaries
Transformer [16] is a deep learning architecture based on multi-head self-attention (MHSA) shown to the better performing than the CNN for a wide variety of tasks. An attention mechanism [17] is a mapping that takes a query, along with key-value pairs from the input sequence to produce an output representation. These input sequences including the queries, keys, values and output sequences, are all represented as matrices. The output is generated by calculating the weighted sum of the input values, with the weights assigned to each value calculated by evaluating the correlation between the query and the respective keys.
@@ II-A Single-head self-attention mechanism
The single-head self-attention (scaled dot-product attention) can be formulated with Q (Query), K (Key) and V (Value) matrices, obtained by linear transformation of the input matrix. The input queries Q, keys K are of dimension and values V of dimension . Q and K are first multiplied and divided by , which is then processed by a softmax operation to construct soft attention before being multiplied with V and returns a weighted value matrix. Finally, V is multiplied with the correlation matrix of Q and K to obtain the self-attention output. This process can be formulated with the following equation.
Where the softmax function converts a matrix of set of real numbers into a probability distribution. For applying softmax function, suppose our input denotes the matrix , where . Softmax function can be computed as followed by .
After doing operation, attention matrix is computed where each row of the attention matrix is equal to 1.
@@ II-B Multi-head self-attention mechanism
Multi-head attention divides the hidden space into multiple subspaces with parameter matrices. And performs the identical computation similar as self-attention mechanism, resulting in multiple Q (Quary), K (Key) and V (Value) matrices. Each of them forms an individual attention that can access information from different subspaces. Multi-head attention linearly projects different Q, K and V matrices h times, where h = number of attention heads. Their results are concatenated and transformed to form the attentive output.
@@ III Problem Formulation
In this section, residential NILM which is the focus of this work has been discussed. A home consists of many appliances, which can be grouped into four categories based on their operational states. Category I: simple on/off appliances like lights and microwaves. Category II: appliances with fixed operating states and repeatable patterns, like fridges and dishwashers. Category III: appliances with variable states, like dimmers and power drills. Category IV: constant-power appliances which operate, like smoke and fire detectors. Category I and Category II appliances are the most common and consume the major portion of electricity in the home, These two appliance categories are the main focus in the existing work, e.g., [18, 10, 11]. The formulation of NILM problem has been given as follows.
@@ IV-B Proposed transformer architecture
Positional information is found to be effective in the literature for transformers to obtain improved performance. A learnable positional vector e is produced and concatenate it with the pooled embedding vector . Formally,
@@
where represent the ground truth and prediction of output sequence divided by maximum power limit. are the appliance state label and prediction. T stands for total time steps that is sequence length and refers to the set of time steps when either the status label is on or the prediction is incorrect. In this equation, we also introduce hyperparameters and for tuning sequence softmax temperature and reduction of absolute error.

### v3

@@ Abstract
Smart grid, NILM, Transformer, Attention.I Introduction
@@ I-A Related works
Deep learning approach: Deep learning has emerged from solving mainstream computer vision problems and influenced other data-driven research fields. Kelly and Knottenbelt are one of the first to propose a deep learning based NILM framework [7] that outperformed classical machine learning methods. The advantage of deep learning architectures is that they can directly process the raw power signal without the need for manual feature extraction or data tagging. They also accommodate noise signals and unknown appliances without assuming that the aggregated power signal fully represents the sum of the exact individual appliance power signal. Zhang et al. [8] enhanced their convolutional neural network (CNN) model’s capacity which surpassed the previous methods. Deep learning research has further improved NILM algorithm’s performance through various strategies, constructing hybrid models using multiple structures such as CNN, long short-term memory (LSTM) and generative adversarial network (GAN) [9, 10]. However, CNN models effectively capture the local features within power consumption sequences but they may not adequately capture global features and dependency correlations between different positions in the sequence.
Attention mechanisms [11], originally formulated in the field of natural language processing, happen to show promising results with NILM problems, especially with transformers. Yue’s BERT4NILM [12] incorporates a bidirectional transformer architecture. Followed by their work methods such as, Sykiotis’s ELECTRIcity-NILM [13], Wang’s Midformer [14], Kamyshev’s COLD,[15] Yue’s ELTransformer[16], Zhou’s TTRNet[17] employed transformer in NILM. Transformer’s capability to capture long-range temporal dependencies made them highly applicable for time-series power consumption data.
@@ II Preliminaries
The Transformer [18] is a deep learning architecture based on multi-head self-attention (MHSA) that has outperformed CNNs across various tasks. The attention mechanism [19] maps a query and key-value pairs from the input to produce an output representation. Queries, keys, values, and outputs are represented as matrices, and the output is computed by calculating the weighted sum of the input values, with weights based on the correlation between the query and keys.
@@ II-A Single-head self-attention mechanism
The single-head self-attention (scaled dot-product attention) can be formulated using the matrices (Query), (Key), and (Value), which are obtained through linear transformations of the input matrix. The queries and keys have dimensions , while the values have dimensions . The product of and is scaled by , then passed through a operation to form soft attention, which is then multiplied by to yield a weighted value matrix. Finally, is multiplied by the correlation matrix of and to compute the self-attention output. This process is represented by the following equation:
Here, the softmax function converts the matrix of real values into a probability distribution. Let represent the matrix , where . The softmax function is applied element-wise across each and can be computed as follows:
After applying the operation, the result is , where each row of the attention matrix sums to 1.
@@ II-B Multi-head self-attention mechanism
Multi-head attention divides the hidden space into multiple subspaces with parameter matrices and performs computations similar to the self-attention mechanism, resulting in multiple (Query), (Key), and (Value) matrices. Each of them forms an individual attention that can access information from different subspaces. Multi-head attention linearly projects different matrices times, where is the number of attention heads. The results are concatenated and transformed to produce the final attention output:
@@ III Problem Formulation
In this section, residential NILM, which is the focus of this work, is discussed. A home consists of many appliances, which can be grouped into four categories based on their operational states:
Category I: Simple on/off appliances such as lights and microwaves. Category II: Appliances with fixed operating states and repeatable patterns, like fridges and dishwashers. Category III: Appliances with variable states, such as dimmers and power drills. Category IV: Constant-power appliances like smoke and fire detectors. Category I and Category II appliances are the most common and consume the majority of electricity in the home. These two appliance categories are the main focus in existing work, e.g., [10, 12, 13]. The formulation of the NILM problem is given as follows:
@@ IV-B Proposed transformer architecture
Positional information is found to be effective in the literature for transformers to obtain improved performance. A learnable positional vector is produced and concatenate it with the pooled embedding vector . Formally,
@@
where represent the ground truth and prediction of the output sequence divided by the maximum power limit. are the appliance state label and prediction. stands for the total time steps, i.e., the sequence length, and refers to the set of time steps when either the status label is on or the prediction is incorrect. In this equation, we also introduce hyperparameters and for tuning the sequence softmax temperature and reduction of absolute error.